# Interpretable Concept Bottlenecks to Align Reinforcement Learning Agents

### v1

@@ Abstract
Reward sparsity, difficult credit assignment, and misalignment are only a few of the many issues that make it difficult, if not impossible, for deep reinforcement learning (RL) agents to learn optimal policies. Unfortunately, the black-box nature of deep networks impedes the inclusion of domain experts who could interpret the model and correct wrong behavior. To this end, we introduce Successive Concept Bottlenecks Agents (SCoBots), which make the whole decision pipeline transparent via the integration of consecutive concept bottleneck layers. SCoBots make use of not only relevant object properties but also of relational concepts. Our experimental results provide strong evidence that SCoBots allow domain experts to efficiently understand and regularize their behavior, resulting in potentially better human-aligned RL. In this way, SCoBots enabled us to identify a misalignment problem in the most simple and iconic video game, Pong, and resolve it.
@@ 1 Introduction
Reinforcement learning (RL) agents, particularly deep RL agents are prone to suffering from a variety of issues that hinder them from learning optimal, generalisable action policies. Prominent examples are reward sparsity (Andrychowicz et al., 2017) and difficult credit assignment (Raposo et al., 2021; Wu et al., 2023). A more recent, yet pressing issue is the goal misalignment problem (di Langosco et al., 2022). It occurs when an agent optimizes a different side-goal that is aligned with the target goal at training time, but not at test time (Koch et al., 2021; di Langosco et al., 2022). We discover that such misalignment can occur in the simplest and most iconic ALE (Mnih et al., 2013) game, Pong, in which an agent’s target goal is to play a ball game against an enemy. However, due to the nature of the automated enemy an agent can learn to focus on the position of the enemy paddle for anticipating where to position its own paddle for catching the ball, rather than the position of the ball itself. This can lead to failure of the agent when the enemy is occluded (cf. Fig. 1 (top)). This phenomenon is more generally described by the umbrella term shortcut learning (Geirhos et al., 2020)) in which a model learns good decision rules for the provided training set that, however, fail to transfer to novel data sets. Importantly, if left unchecked, it may lead to a lack of model generalization and dangerous and unintuitive failures when such models are deployed in real-world settings.
@@
Figure 1: Successive Concept Bottlenecks Agents (SCoBots) allow for easy inspection and revision. Top: Deep RL agents trained on Pong produce high playing scores ( Refer to caption ) while importance map explanations suggest sensible underlying reasons of taking an action. However, when the enemy is hidden, the deep RL agent fails to even catch the ball ( Refer to caption ) without clear reasons. Bottom: SCoBots, on the other hand, allow for multi-level inspection of the reasoning behind the action selection , e.g., at a relational concept, but also action level. Moreover, they allow for users to easily intervene on them ( Refer to caption ) to prevent the agents from focusing on potentially misleading concepts. In this way SCoBots can mitigate RL specific caveats like goal misalignment.
@@
A recently popular approach for identifying such shortcut behaviour is via methods from the field of eXplainable AI (XAI) (e.g. (Schramowski et al., 2020; Lapuschkin et al., 2019)) where the goal of XAI methods is to identify the reasons behind a model’s decisions (Guidotti et al., 2019; Ras et al., 2022; Roy et al., 2022; Saeed & Omlin, 2023). Many methods of XAI however suffer from issues such as faithfulness (the explanation does not faithfully present a model’s underlying decision process)(Hooker et al., 2019; Chan et al., 2022; DeYoung et al., 2020) and low-level semantics (explanations are made at the level of the input data e.g. pixels and not the relevant concepts that are extracted from these pixels)(Stammer et al., 2021; Kambhampati et al., 2021; Teso et al., 2023). A recent branch of research therefore focuses on developing models that provide inherent, concept-based explanations. Prominent examples of this idea are concept bottlenecks models (CBMs) which, among other things, have been shown to provide predictive performances on par with standard deep learning approaches for supervised image classification (Koh et al., 2020). More importantly, CBMs allow to identify and effectively revise incorrect model behaviour on a higher, concept level (Stammer et al., 2021).
@@
Despite the success and advantageous properties of CBMs up to now, no research has been performed on their integration into RL, particularly for tackling RL-specific issues discussed above. A reason for this is that previous concept-based models depend only on concepts that are at a single object level, e.g., object color, shape etc… However, objects in RL environments such as ALE games interact with another (Kaiser et al., 2019). Therefore, in order to adequately solve such tasks additional relational concepts are necessary that, however, build upon these single object concepts.
@@
For this purpose, in this work we introduce Successive Concept Bottleneck Agents (SCoBots). SCoBots (cf. Fig. 1 (bottom)) hereby integrate multiple, successive concept bottlenecks into their decision processes where each bottleneck layer consists of concepts representations that integrate the concept representations of the previous bottleneck layer(s). Importantly, given a set of predefined concept functions that automatically set objects and concepts of initial bottleneck layers in relation to another the final bottleneck layer consists of valuable relational concept representations. SCOBots thus represent inherently explainable RL agents which in comparison to deep RL agents allow for inspecting and revising their learned decision policies at multiple levels of their reasoning processes, from single object concepts, to relational concepts to the actual action decision.
@@
In our evaluations on benchmark Atari games we provide experimental evidence not just of the performance abilities of SCoBots that is on par with deep RL agents. More importantly we showcase their ability to provide valuable explanations and their ability to mitigate a variety of RL specific issues, from reward sparsity to misalignment problems. Importantly, in the course of our evaluations we identify previously unknown shortcut behaviour of deep RL agents even on the arguably simplest of all Atari games, Pong. Utilizing the advantages of SCoBots this behaviour is, however, easily removed. Ultimately our work thereby illustrates the severity of goal misalignment issues in RL and the importance of being able to mitigate these and other RL specific issues via relational concept based models.
@@
We introduce Successive Concept Bottleneck agents (SCoBots) that integrate relational concept representations into their predictions.
We show that SCoBots allow to follow their internal decision processes.
@@
We show that they allow human interactions that lead to mitigating different RL specific learning issues.
We proceed as follows. We first formally introduce SCoBots and discuss their specific properties. We continue with extensive experimental evaluations on Atari game environments via several analysis and metrics. Before concluding, we touch upon related work.
@@ 2 Successive Concept Bottleneck Agents
Let us start with providing the background notations on RL via markov decision processes (MDPs) as well as concept bottleneck models (CBMs). In this work, we represent an RL problem through the framework of a Markov Decision Process, $\mathcal{M}=<\mathcal{S},\mathcal{A},P_{s,a},R_{s,a}>$, with $\mathcal{S}$ as the state space, $\mathcal{A}$ the set of available actions, $P(s,a)$ the transition probability, and $R(s,a)$ the immediate reward function, obtained from the environment. Classic deep RL policies, $\pi_{\theta}(s)=P(A=a|S=s)$ with parameters $\theta$, are usually black-box models that process raw input states, e.g. a set of frames, and provide a distribution over the action space (Mnih et al., 2015; van Hasselt et al., 2016).
Concept bottleneck models (CBMs) initiate a learning process by extracting relevant concepts from raw input e.g. image data. These concepts can e.g. represent the color or position of a an Atari character. Subsequently, these extracted concepts are used for downstream tasks such as classification tasks. Formally, a bottleneck model $g:x\rightarrow c$ transforms an input $x\in\mathcal{X}$ into a concept representation c. Next, the predictor network $f:c\rightarrow y$ uses this representation to generate the final target output $y\in\mathcal{Y}$.
The main body of research on CBMs focuses on image classification tasks, where extracted concepts represent attributes of objects. In contrast, RL agents must learn not just from static, labeled data but also through interaction with dynamic environments. RL tasks thus often necessitate relational reasoning, as they involve understanding the relationships between depicted instances that evolve through time and interact with one another. This is crucial for learning effective policies in complex, dynamic environments. In this manuscript, we use specific fonts to distinguish object properties and relations.
@@ 2.1 Building inspectable ScoBots
An underlying assumption of SCoBots is that their processing steps should be inspectable and understandable by a human user. This stands in contrast e.g. to unsupervised CBM approaches (Jabri et al., 2019; Zhou et al., 2019; Srinivas et al., 2020) where there is no guarantee for learning human-relatable concepts. SCoBots rather take a different approach by dividing the concept learning and RL-specific action selection into several inspectable steps that are grounded in human-understandable concept representations. These steps are described herafter and depicted in Fig. 2.
@@
Similar to other RL agents, SCoBots process the last $n$ frames $\{x_{t}\}_{t=T-n}^{T}$, observed from the environment. For each frame, SCoBots employ an initial concept extraction method as is standard in other works on CBMs (Stammer et al., 2021; Koh et al., 2020) to extract objects and their properties. Specifically, we start from an initial bottleneck model, $\omega_{\theta_{1}}(\cdot)$, that is parameterized by parameter set, $\theta_{1}$. Given an image, $x_{t}$, this model provides a set of $c_{t}$ object representations, $\omega_{\theta_{1}}(x_{t})=\Omega_{t}=\{o_{t}^{j}\}_{j=1}^{c_{t}}$, where each object representation corresponds to a tensor of different extracted properties of that object, (e.g. the object category, position (i.e. its $x,y$ coordinates), size, … etc). As in previous works on CBMs this bottleneck model, $\omega_{\theta_{1}}(\cdot)$, can correspond to a model that was supervisedly pretrained for predicting these object properties from an image.
@@
One of the major differences to previous work on CBMs is that SCoBots now extract and utilize relational concepts that are based on the previously extracted objects and their properties. Formally, SCoBots utilize a successive bottleneck model, $\mu_{\mathcal{F}}(\cdot)$, called the relation extractor. This model, $\mu$, is parameterized by a predetermined set of relational functions, $\mathcal{F}$, that are used to extract a set of $d_{t}$ relational concepts. These relations are based on each individual object or property for unary relations or combinations of these for n-ary relations and are denoted as $\mu_{\mathcal{F}}(\Omega_{t})=\Gamma_{t}=\{g_{t}^{k}\}_{k=1}^{d_{t}}$. Without strong prior knowledge $\mathcal{F}$ can initially correspond to universal object relations such as distance and speed. However, this set can easily be updated on the fly by a human user e.g. with additional, novel relation functions.
@@
Finally, SCoBots employ an action selector, $\rho_{\theta_{2}}$, on the relational concepts to select the best action, $a_{t}$, given the initial state, $s_{t}$ (e.g. image $x_{t}$) and that is parameterized by $\theta_{2}$}. Up to now all extracted concepts in SCoBots, both $\Omega_{t}$ and $\Gamma_{t}$, represent human-understandable concept representations. To guarantee understandability also within the action selection step of SCoBots we require to embed a specifically interpretable action selector. While neural networks, a standard choice in RL literature, are performative and easy to train using differentiable optimization methods they lack this required interpretability feature. However, (Çağlar Aytekin, 2022) have recently shown the equivalence between ReLU-based neural networks and decision trees, which in contrast represent inherently interpretable models. Thus, to trade-off the issues of flexibility and performance vs interpretability SCoBots thus break down the action selection process by initially training a small ReLU-based neural network action selector, $\widetilde{\rho}_{\theta^{\prime}_{2}}$, via gradient-based RL optimization. After this, $\widetilde{\rho}_{\theta^{\prime}_{2}}$ is finally distilled into a decision tree $\rho_{\theta_{2}}$. Lastly, note that one can add a residual link from the initial object concepts, $\Omega_{t}$, to the relational concepts, $\Gamma_{t}$ such that the action selector can, if necessary, also make decisions based on basic object properties, e.g., the height of an object.
@@
where $\theta=(\theta_{1},\theta_{2},\mathcal{F}$) are the policy parameters. For simplicity, we will discard the parameter notations in the rest of the manuscript. Further explanations of the input and output space of each module, as well as the properties and relational functions used in this work are provided in App.A.6 and App. A.7.
Instead of jointly learning object detection, concept extraction, and policy search, SCoBots enable the optimization of each policy component independently. This separation of training for different components (similar to that of (Koh et al., 2020)) particularly reduces the complexity of the overall optimization problem. In this way, the object-concept extractor, $\omega$, can be optimized by supervised (Redmon et al., 2016) or unsupervised (Lin et al., 2020; Delfosse et al., 2023c) object detection methods. On the other hand, the set of suitable relations, $\mathcal{F}$, as well which objects to apply them on, may be adapted via evolutionary algorithms (Peng et al., 2018; Yang et al., 2022), supervision from large language models (Wu et al., 2023), via prior task-knowledge (Reid et al., 2022), or neural guidance from a pretrained fully deep agent (Delfosse et al., 2023b). Conclusively, the action selector, $\rho$ can be optimized based on the direct RL signal.

### v2

@@ Abstract
Goal misalignment, reward sparsity and difficult credit assignment are only a few of the many issues that make it difficult for deep reinforcement learning (RL) agents to learn optimal policies. Unfortunately, the black-box nature of deep neural networks impedes the inclusion of domain experts for inspecting the model and revising suboptimal policies. To this end, we introduce Successive Concept Bottleneck Agents (SCoBots), that integrate consecutive concept bottleneck (CB) layers. In contrast to current CB models, SCoBots do not just represent concepts as properties of individual objects, but also as relations between objects which is crucial for many RL tasks. Our experimental results [^1] Code available at https://github.com/k4ntz/SCoBots provide evidence of SCoBots’ competitive performances, but also of their potential for domain experts to understand and regularize their behavior. Among other things, SCoBots enabled us to identify a previously unknown misalignment problem in the iconic video game, Pong, and resolve it. Overall, SCoBots thus result in more human-aligned RL agents.
@@ 1 Introduction
Reinforcement learning (RL) agents, particularly deep RL agents are prone to suffer from a variety of issues that hinder them from learning optimal or generalisable policies. Prominent examples are reward sparsity (Andrychowicz et al., 2017) and difficult credit assignment (Raposo et al., 2021; Wu et al., 2023). A more recent, yet pressing issue is the goal misalignment problem. It occurs when an agent optimizes a different side-goal, that is aligned with the original target goal during training (Koch et al., 2021), but not at test time. Such misalignments can be difficult to identify (di Langosco et al., 2022). For instance in this work we discover for the first time that such a misalignment can occur in the one of the oldest and most iconic Atari games, Pong (cf. Fig. 1, top). In Pong, the agent’s target goal is to catch a ball with its own paddle and to return it passed the enemy’s one (Mnih et al., 2013). Due to the automated enemy, that constantly follows the ball, agents can learn to focus on the position of the enemy paddle’s for placing their own paddle, rather than the position of the ball itself. This phenomenon is more generally described by the umbrella term shortcut learning (Geirhos et al., 2020) in which a model learns decision rules for the provided training set, that fail to transfer to novel data sets. If left unchecked, it may lead to a lack of model generalization and dangerous and unintuitive failures when such models are deployed in real-world settings (Advani & Saxe, 2017; Zhang et al., 2021).
@@
Figure 1: Successive Concept Bottlenecks Agents (SCoBots) allow for easy inspection and revision. Top: Deep RL agents trained on Pong produce high playing scores with importance map explanations that suggest sensible underlying reasons of taking an action ( Refer to caption ). However, when the enemy is hidden, the deep RL agent fails to even catch the ball without clear reasons ( Refer to caption ). Bottom: SCoBots, on the other hand, allow for multi-level inspection of the reasoning behind the action selection , e.g., at a relational concept, but also action level. Moreover, they allow for users to easily intervene on them ( Refer to caption ) to prevent the agents from focusing on potentially misleading concepts. In this way SCoBots can mitigate RL specific caveats like goal misalignment.
@@
A recently popular approach for identifying such shortcut behaviour is via methods from the field of eXplainable AI (XAI), whose aim is to identify the reasons behind a model’s decisions (Lapuschkin et al., 2019; Guidotti et al., 2019; Schramowski et al., 2020; Ras et al., 2022; Roy et al., 2022; Saeed & Omlin, 2023). Many methods of XAI however suffer from issues such as a lack of faithfulness, i.e. the explanation does not faithfully present the model’s underlying decision process (Hooker et al., 2019; Chan et al., 2022; DeYoung et al., 2020), and coarse, low-level explanation semantics. For example, importance-map explanations indicate the importance of an input element for a decision without indicating why this element is important (Kambhampati et al., 2021; Stammer et al., 2021; Teso et al., 2023). A recent branch of research therefore focuses on developing models that provide inherent concept-based explanations. Prominent examples of this idea are concept bottlenecks models (CBMs) which have been shown to provide predictive performances on par with standard deep learning approaches for supervised image classification (Koh et al., 2020). More importantly, CBMs allow to identify and effectively revise incorrect model behaviour on a higher, concept level (Stammer et al., 2021).
@@
Despite the success and advantageous properties of CBMs for tackling RL-specific issues discussed above, no research has however been performed for their integration into RL. A reason for this is that previous concept-based models depend only on concepts that are at a single object level, e.g., object color, shape, etc. However, objects in RL environments, even as simple as Atari games, interact with one another (Kaiser et al., 2019). In order to adequately solve such RL tasks via explicit concept representations, agents thus need to also extract necessary relational concepts.
@@
In this work we therefore introduce Successive Concept Bottleneck Agents (SCoBots, cf. Fig. 1, bottom) as an important extension of the original concept bottleneck approach to RL. SCoBots integrate successive concept bottlenecks into their decision processes where each bottleneck layer provides concept representations that integrate the concept representations of the previous bottleneck layer. Specifically, provided a set of predefined concept functions SCoBots automatically extract relational concept representations based on objects and their properties of the initial bottleneck layers. Finally, the set of relational and object concept representations are used for optimal action selection. SCoBots thus represent inherently explainable RL agents and in comparison to deep RL agents, they allow for inspecting and revising their learned decision policies at multiple levels of their reasoning processes: from the single object properties, through the relational concepts to the actual action decision.
@@
In our evaluations on the iconic Atari Learning Environments (Mnih et al., 2013) (ALE), we provide experimental evidence of the performance abilities of SCoBots that is on par with deep RL agents. More importantly, we showcase their ability to provide valuable explanations and the potential of mitigating a variety of RL specific issues, from reward sparsity to misalignment problems, via simple guidance from domain experts. In the course of our evaluations we identify previously unknown shortcut behaviour of deep RL agents, even on a simple game as Pong. By utilizing the interaction capabilities of SCoBots, this behaviour is easily removed. Ultimately, our work thereby illustrates the severity of goal misalignment issues in RL and the importance of being able to mitigate these and other RL specific issues via relational concept based models.
@@
We introduce Successive Concept Bottleneck agents (SCoBots) that integrate relational concept representations into their decision processes.
We show that SCoBots allow to inspect their internal decision processes.
@@
We show that they allow for human interactions for mitigating various RL specific learning issues.
We proceed as follows. We first formally introduce SCoBots and discuss their specific properties. We continue with experimental evaluations via several metrics and analysis. Before concluding, we touch upon related work.
@@ 2 Successive Concept Bottleneck Agents
Let us start with providing the background notations of RL problems, modelled as markov decision processes (MDPs), and also of concept bottleneck models (CBMs). In this work, we represent an RL problem through the framework of a Markov Decision Process, $\mathcal{M}=<\mathcal{S},\mathcal{A},P_{s,a},R_{s,a}>$, with $\mathcal{S}$ as the state space, $\mathcal{A}$ the set of available actions, $P(s,a)$ the transition probability, and $R(s,a)$ the immediate reward function, obtained from the environment. Classic deep RL policies, $\pi_{\theta}(s)=P(A=a|S=s)$ parametrized by $\theta$, are usually black-box models that process raw input states, e.g. a set of frames, to provide a distribution over the action space (Mnih et al., 2015; van Hasselt et al., 2016).
Concept bottleneck models (CBMs) initiate a learning process by extracting relevant concepts from raw input (e.g. image data). These concepts can represent the color or position of a depicted object. Subsequently, these extracted concepts are used for downstream tasks such as image classification. Formally, a bottleneck model $g:x\rightarrow c$ transforms an input $x\in\mathbb{R}^{D}$ with $D$ dimensions into a concept representation $c\in\mathbb{R}^{k}$ (a vector of $k$ concepts). Next, the predictor network $f:c\rightarrow y$ uses this representation to generate the final target output (e.g. $y\in\mathbb{R}$ for classification problems).
The main body of research on CBMs focuses on image classification tasks, where extracted concepts represent attributes of objects. In contrast, RL agents must learn not just from static data but also through interaction with dynamic environments. RL tasks thus often require relational reasoning, as they involve understanding the relationships between instances that evolve through time and interact with another. This is crucial for learning effective policies in complex, dynamic environments. Note that, in the following descriptions, we use specific fonts to distinguish objects’ properties from their relations.
@@ 2.1 Building inspectable ScoBots
An underlying assumption of SCoBots is that their processing steps should be inspectable and understandable by a human user. This stands in contrast to e.g. unsupervised CBM approaches (Jabri et al., 2019; Zhou et al., 2019; Srinivas et al., 2020), where there is no guarantee for learning human-aligned concepts. SCoBots rather take a different approach by dividing the concept learning and RL-specific action selection into several inspectable steps that are grounded in human-understandable concept representations. These steps are described hereafter and depicted in Fig. 2.
@@
Similar to other RL agents, SCoBots process the last $n$ observed frames from a sequence of images, $s_{t}=\{x_{i}\}_{i=t-n}^{t}$. For each frame, SCoBots need an initial concept extraction method to extract objects and their properties, as in previous works on CBMs (Stammer et al., 2021; Koh et al., 2020). Specifically, we start from an initial bottleneck model, $\omega_{\theta_{1}}(\cdot)$, that is parameterized by parameter set, $\theta_{1}$. Given a state, $s_{t}$, this model provides a set of $c_{i}$ object representations per frame, $\omega_{\theta_{1}}(s_{t})=\Omega_{t}=\{\{o_{i}^{j}\}_{j=1}^{c_{i}}\}_{i=t-n}^{t}$, where each object representation corresponds to a tensor of different extracted properties of that object, (e.g. its category, position (i.e. $x,y$ coordinates), etc.). As done in previous works on CBMs, the bottleneck model of our SCoBots, $\omega_{\theta_{1}}$, can correspond to a model that was supervisedly pretrained for extracting the objects and their properties from images.
@@
One of the major differences to previous work on CBMs is that SCoBots further extract and utilize relational concepts that are based on the previously extracted objects and their properties. Formally, SCoBots utilize a consecutive bottleneck model, $\mu_{\mathcal{F}}(\cdot)$, called the relation extractor. This model, $\mu$, is parameterized by a predetermined set of transparent relational functions, $\mathcal{F}$, used to extract a set of $d_{t}$ relational concepts. These relations are based on each individual object (and its properties) for unary relations or on combinations of objects for n-ary relations and are denoted as $\mu_{\mathcal{F}}(\Omega_{t})=\Gamma_{t}=\{g_{t}^{k}\}_{k=1}^{d_{t}}$. Without strong prior knowledge, $\mathcal{F}$ can initially correspond to universal object relations such as distance and speed. However, this set can easily be updated on the fly by a human user with e.g. additional, novel relational functions. Note that $\mathcal{F}$ can include the identity, to let the relation extractor pass through initial object concepts from $\Omega_{t}$ to the relational concepts $\Gamma_{t}$.
@@
Finally, SCoBots employ an action selector, $\rho_{\theta_{2}}$, parameterized by $\theta_{2}$}, on the relational concepts to select the best action, $a_{t}$, given the initial state, $s_{t}$ (i.e. the set of frames). Up to now, all extracted concepts in SCoBots, both $\Omega_{t}$ and $\Gamma_{t}$, represent human-understandable concept representations. To guarantee understandability also within the action selection step of SCoBots, we need to embed an interpretable action selector. While neural networks, a standard choice in RL literature, are performative and easy to train using differentiable optimization methods, they lack this required interpretability feature. However, (Çağlar Aytekin, 2022) have recently shown the equivalence between ReLU-based neural networks and decision trees, which in contrast represent inherently interpretable models. To trade-off these issues of flexibility and performance vs interpretability, SCoBots thus break down the action selection process by initially training a small ReLU-based neural network action selector, $\widetilde{\rho}_{\theta^{\prime}_{2}}$, via gradient-based RL optimization. After this, $\widetilde{\rho}_{\theta^{\prime}_{2}}$ is finally distilled into a decision tree $\rho_{\theta_{2}}$. Lastly, note that one can add a residual link from the initial object concepts, $\Omega_{t}$, to the relational concepts, $\Gamma_{t}$ such that the action selector can, if necessary, also make decisions based on basic object properties, e.g., the height of an object.
@@
where $\theta=(\theta_{1},\theta_{2},\mathcal{F}$) constitutes the set of policy parameters. For simplicity, we will discard the parameter notations in the rest of the manuscript. Further explanations of the input and output space of each module, as well as the properties and relational functions used in this work are provided in the appendix (cf. App.A.5 and App. A.6).
Instead of jointly learning object detection, concept extraction, and policy search, SCoBots enable the optimization of each policy component independently. Separating the training procedure of different components (similar to that of (Koh et al., 2020)) particularly reduces the complexity of the overall optimization problem.